Which of the following provides the greatest annual return? 16% compounded quarterly 15% compounded continuously 16% compounded annually all of these are exactly equal 15.2% compounded daily

*Please show math step by step, thumbs up guaranteed <3
Answer #1

Calculating Effective Annual Rate(EAR) for each of the option to determine the greatest annual return:-

i) 16% compounded Quarterly

EAR = (1+\frac{r}{m})^m-1

where, r = Interest rate = 16%

m = No of times compounding in a year = 4

EAR = (1+\frac{0.16}{4})^4-1

EAR = 1.169859 - 1

EAR = 16.9859%

ii) 15% compounded continously

EAR = e^r-1

where, r = Interest rate = 15%

EAR =e^{0.15}-1

EAR = 1.161834 - 1

EAR = 16.1834%

iii) 16% compounded annually

EAR = (1+\frac{r}{m})^m-1

where, r = Interest rate = 16%

m = No of times compounding in a year = 1

EAR =(1+\frac{0.16}{1})^1-1

EAR = 1.16 - 1

EAR = 16%

iv) 15.20% compounded Daily

EAR = (1+\frac{r}{m})^m-1

where, r = Interest rate = 15.20%

m = No of times compounding in a year = 365

EAR =(1+\frac{0.1520}{365})^{365}-1

EAR = 1.164123 - 1

EAR = 16.4123%

As EAR of 16% componded Quarterly is highest among the other Options, it is better to choose "16% Compounded Quarterly".

If you need any clarification, you can ask in comments.     

If you like my answer, then please up-vote as it will be motivating